# SVTRv2: CTC Beats Encoder-Decoder Models in Scene Text Recognition

### v1

@@ 3.5 Optimization Objective
During training, The optimization objective is to minimize the loss , which comprises and , as listed below:
@@ 4.1 Datasets and Implementation Details
We evaluate SVTRv2 across multiple benchmarks covering diverse scenarios. They are: 1) six common regular and irregular benchmarks (Com), including ICDAR 2013 (IC13) [27], Street View Text (SVT) [45], IIIT5K-Words (IIIT5K) [33], ICDAR 2015 (IC15) [26], Street View Text-Perspective (SVTP) [35] and CUTE80 [1]. For IC13 and IC15, we use the versions with 857 and 1811 images, respectively; 2) the test set of the recent Union14M-L benchmark (U14M) [25], which includes seven challenging subsets: Curve, Multi-Oriented (MO), Artistic, Contextless (Cless), Salient, Multi-Words (MW) and General; 3) occluded scene text dataset (OST) [48], which is categorized into two subsets based on the degree of occlusion: weak occlusion (OSTw) and heavy occlusion (OSTh); 4) long text benchmark (LTB) [13], which includes 3376 samples of text length from 25 to 35; 5) the test set of BCTR [6], a Chinese text recognition benchmark with four subsets: Scene, Web, Document (Doc) and Hand-Writing (HW).
For English recognition, we train models on real-world datasets, from which the models exhibit stronger recognition capability [4, 25, 37]. There are three large-scale real-world training sets, i.e., the Real dataset [4], REBU-Syn [37], and the training set of Union14M-L (U14M-Train) [25]. However, they all overlap with U14M (detailed in Sec. 8 in Supplementary) across the seven subsets, leading to data leakage, which makes them unsuitable for training models. To resolve this, we introduce a filtered version of Union14M-L training set, termed as U14M-Filter, by filtering these overlapping instances. This new dataset is used to train SVTRv2 and 24 mainstream methods we reproduced.
For Chinese recognition, we train models on the training set of BCTR [6]. Unlike previous methods that train separately for each subset, we trained the model on an integrated dataset and then evaluated it on the four subsets.
We use AdamW optimizer [30] with a weight decay of 0.05 for training. The LR is set to and batchsize is set to 1024. One cycle LR scheduler [43] with 1.5/4.5 epochs linear warm-up is used in all the 20/100 epochs, where a/b means a for English and b for Chinese. For English model, we take SVTRv2 without SGM as the pre-trained and fine-tuned SVTRv2 with SGM with the same above settings. Word accuracy is used as the evaluation metric. Data augmentation like rotation, perspective distortion, motion blur and gaussian noise, are randomly performed and the maximum text length is set to 25 during training. The size of the character set is set to 94 for English and 6624 [28] for Chinese. In experiments, SVTRv2 means SVTRv2-B unless specified. All models are trained with mixed-precision on 4 RTX 4090 GPUs.
@@ 4.2 Ablation Study
Effectiveness of MSR. We group the Curve and MO text in U14M based on the aspect ratio Ri. As shown in Tab. 1, the majority of irregular texts fall within R1 and R2, where they are particularly prone to distortion when resized to a fixed size (see Fixed32×128 in Fig. 4). In contrast, MSR demonstrates significant improvements of 15.3% in R1 and 5.2% in R2 compared to Fixed32×128. Meanwhile, a large fixed-size Fixed64×256, although improving the accuracy compared to the baseline, still performs worse than our MSR by clear margins. The results strongly confirm our hypothesis that undesired resizing would hurt the recognition. Our MSR effectively mitigates this issue, providing better visual features thus enhancing the recognition accuracy.
Effectiveness of FRM. We ablate the two rearrangement sub-modules (Horizontal (H) rearranging and Vertical (V) rearranging). As shown in Tab. 1, compared to without FRM (w/o FRM), they individually improve accuracy by 2.03% and 0.71% on MO, and they together result in a 2.46% gain. Additionally, we explore using a Transformer block (+ TF1) to learn the rearrangement matrix holistically, whose effectiveness is less obvious. The most probable reason is that this scheme does not well distinguish between vertical and horizontal orientations. In contrast, FRM performs feature rearrangement in both directions, making it highly sensitive to text irregularity, and thus facilitating accurate CTC alignment. As shown in the left five cases in Fig. 4, FRM successfully recognizes reverse instances, providing strong evidence of FRM’s effectiveness.
@@
| Method | Venue | Encoder | Common Benchmarks | Avg | Union14M Benchmarks | Avg | LTB | OST | Size | FPS | ||||||||||||||||||||||||||
@@ 4.3 Comparison with State-of-the-arts
To demonstrate the effectiveness of SVTRv2 in English, we compare it with 24 popular STR methods. All the models are tested on the newly constructed U14M and the results are given in Tab. 3. SVTRv2-B ranks the top in 12 of the 15 evaluated scenarios, It almost outperforms all the EDTRs in every scenario, showing a clear accuracy advantage. Meanwhile, it still enjoys a small model size and a significant speed advantage. Specifically, compared to MAERec, the best-performed existing model on U14M, SVTRv2-B shows an accuracy improvement of 0.97% and 8 faster inference speed. Compared to CPPD, which is known for its accuracy-speed tradeoff, SVTR-B runs faster than 10%, along with a 4.23% accuracy increase on U14M. Regarding OST, as illustrated in the right part of Fig. 4, SVTR-B relies solely on a single visual model but achieves comparable accuracy to PARSeq, which employed the advanced permuted language model and is the best-performed existing model on OST. In the case of long text recognition, where a large portion of EDTRs are incapable of recognizing, SVTR-B outperforms LISTER, the best EDTR method on LTB, by 13%, demonstrating the remarkable scalability of SVTRv2. In addition, SVTRv2-T and SVTRv2-S, the two smaller models also show leading accuracy compared with models of similar sizes, offering solutions with different accuracy-speed tradeoff.
Two observations are derived when looking into the results on Curve and MO. First, SVTRv2 models significantly surpass existing CTC models. For example, compared to SVTR-B, SVTRv2-B gains prominent accuracy improvements of 14.4% and 44.5%, respectively. Second, as shown in Tab. 4, comparing with previous methods employing the rectification modules [40, 36, 59, 62, 54, 11, 61] and the attention-based decoder [38, 31, 5, 51, 56, 25, 54, 52, 47, 29] to recognize irregular text, SVTRv2 also performs better than these methods on Curve and MO. In Fig. 4, the rectification module (TPS) and the attention-based decoder (MAERec*) do not recognize the extremely curved and rotated text correctly, in contrast, SMTR successes. Moreover, as demonstrated by the results on LTB in Tab. 4 and Fig. 5, TPS and MAERec* both do not effectively recognize long text, while SVTRv2 circumvents this limitation. These results indicate that our proposed modules successfully address the challenge of handling irregular text that existing CTC models encountered, while preserving CTC’s proficiency in recognizing long text.
SVTRv2 models also exhibit strong performance in Chinese text recognition (see Tab. 5), where SVTRv2-B achieve state of the art. The result underscores its great adaptability to different languages. To sum, we evaluate SVTRv2 across a wide range of scenarios. The results consistently confirm that this CTC model beats leading EDTRs.
| R1 | R2 | R3 | R4 | Curve | MO | Com | U14M | LTB | ||
@@ 5 Conclusion
In this paper, we have presented SVTRv2, an accurate and efficient CTC-based STR method. SVTRv2 is featured by developing the MSR and FRM modules to tackle the text irregular challenge, and devising the SGM module to endow linguistic context to the visual model. These upgrades maintain the simple architecture of CTC models, thus they remain quite efficient. More importantly, our thorough validation on multiple benchmarks demonstrates the effectiveness of SVTRv2. It achieves leading accuracy and inference speed in various challenging scenarios covering regular, irregular, occluded, Chinese and long text, convincingly indicating that SVTRv2 has beat EDTRs in scene text recognition. In addition, we retrain 24 methods from scratch on U14M-Filter without data leakage, constituting a comprehensive and reliable benchmark. We hope that SVTRv2 and this benchmark will further advance the development of the OCR community.
@@
| ID | Method | Common Benchmarks | Avg | Union14M Benchmarks | Avg | LTB | OST | Size | FPS | |||||||||||||||||||||||||
@@
6 More detail of Ablation Study
SVTRv2 builds upon the foundation of SVTR by introducing several innovative strategies aimed at addressing challenges in recognizing irregular text and modeling linguistic context. The key advancements and their impact are systematically detailed below:
Removal of the Rectification Module and Introduction of MSR and FRM. In the original SVTR, a rectification module is employed to recognize irregular texts. However, this approach negatively impacts the recognition of long texts. To overcome this limitation, SVTRv2 removes the rectification module entirely. To effectively handle irregular text without compromising the CTC model’s ability to generalize to long text, MSR and FRM are introduced.
@@ 6 More detail of Ablation Study
Refinement of Local Mixing Mechanisms. SVTR employs a hierarchical vision transformer structure, leveraging two mixing strategies: Local Mixing is implemented through a sliding window-based local attention mechanism, and Global Mixing employs the standard global multi-head self-attention mechanism. SVTRv2 retains the hierarchical vision transformer structure and the global multi-head self-attention mechanism for Global Mixing. For Local Mixing, SVTRv2 introduces a pivotal change. Specifically, the sliding window-based local attention is replaced with two consecutive group convolutions (Conv2). It is important to highlight that unlike previous CNNs [41, 20, 22], there is no normalization or activation layer between the two convolutions.
Semantic Guidance Module (SGM). The original SVTR model relies solely on CTC framework for both training and inference. However, CTC is inherently limited in its ability to model linguistic context. SVTRv2 addresses this by introducing a Semantic Guidance Module (SGM) during training. SGM facilitates the visual encoder in capturing linguistic information, enriching the feature representation. Importantly, SGM is discarded during inference, ensuring that the efficiency of CTC-based decoding remains unaffected while still benefiting from its contributions during the training phase.
@@ 6.1 Progressive Ablation Experiments
To comprehensively evaluate the contributions of the innovations in SVTRv2, a series of progressive ablation experiments are conducted. Tab 6 outlines the results, with the following observations:
@@
2. Rectification Module Removal (ID1) reveals that while the rectification module (e.g., TPS) improves irregular text recognition accuracy, it hinders the model’s ability to recognize long texts. This confirms its limitations in balancing these tasks.
@@
4. Replacement of Local Attention with Conv2 (ID3): Replacing the sliding window-based local attention with two consecutive group convolutions (Conv2) yields improvements in artistic text, with a 3.0% increase in accuracy. This result highlights the efficacy of convolution-based approaches in capturing character-level nuances, such as strokes and textures, thereby improving its ability to recognize artistic and irregular text styles.
@@
6. Integration of SGM (ID6): Adding SGM yields significant gains on multiple datasets, improving accuracy on OST by 5.11% and Union14M-Benchmark by 2.28%.
It can be summarized as that, by integrating Conv2, MSR, FRM, and SGM, SVTRv2 significantly improves performance in recognizing irregular text and modeling linguistic context over SVTR, while maintaining robust long-text recognition capabilities and preserving the efficiency of CTC-based inference.
@@ 7 SVTRv2 Variants
denotes that the first m mixing blocks in SVTRv2 utilize local mixing, while the last n mixing blocks employ global mixing. Specifically, in SVTRv2-T and SVTR-S, all blocks in the first stage and the first three blocks in the second stage use local mixing. The last three blocks in the second stage, as well as all blocks in the third stage, are global mixing. In the case of SVTRv2-B, all blocks in the first stage and the first two blocks in the second stage use local mixing, whereas the last four blocks in the second stage and all blocks in the third stage adopt global mixing.
@@
| Curve | MO | Artistic | Cless | Salient | MW | General | |
@@
| Method | Venue | Encoder | Common Benchmarks | Avg | Union14M Benchmarks | Avg | Size | |||||||||||||||||||||||||

### v2

@@ 3.5 Optimization Objective
During training, the optimization objective is to minimize the loss , which comprises and as listed below:
@@ 4.1 Datasets and Implementation Details
We evaluate SVTRv2 across multiple benchmarks covering diverse scenarios. They are: 1) six common regular and irregular benchmarks (Com), including ICDAR 2013 (IC13) [27], Street View Text (SVT) [44], IIIT5K-Words (IIIT5K) [33], ICDAR 2015 (IC15) [26], Street View Text-Perspective (SVTP) [35] and CUTE80 [1]. For IC13 and IC15, we use the versions with 857 and 1811 images, respectively; 2) the recent Union14M-Benchmark (U14M) [25], which includes seven challenging subsets: Curve, Multi-Oriented (MO), Artistic, Contextless, Salient, Multi-Words and General; 3) occluded scene text dataset (OST) [47], which is categorized into two subsets based on the degree of occlusion: weak occlusion (OSTw) and heavy occlusion (OSTh); 4) long text benchmark (LTB) [12], which includes 3376 samples of text length from 25 to 35; 5) the test set of BCTR [6], a Chinese text recognition benchmark with four subsets: Scene, Web, Document (Doc) and Hand-Writing (HW).
For English recognition, there are three large-scale real-world training sets, i.e., the Real dataset [4], REBU-Syn [37], and Union14M-L [25]. However, they all overlap with U14M (detailed in Suppl. Sec. 8) across the seven subsets, leading to data leakage, which makes them unsuitable for training models. To resolve this, we introduce a filtered version of Union14M-L, termed as U14M-Filter, by filtering these overlapping instances. This new dataset is used to train SVTRv2 and 24 popular STR methods.
For Chinese recognition, we train models on the training set of BCTR [6]. Unlike previous methods that train separately for each subset, we trained the model on their integration and then evaluated it on the four subsets.
We use AdamW optimizer [30] with a weight decay of 0.05 for training. The LR is set to and batchsize is set to 1024. One cycle LR scheduler [42] with 1.5/4.5 epochs linear warm-up is used in all the 20/100 epochs, where a/b means a for English and b for Chinese. For English models, the training is conducted in two phases: firstly without SGM and then with SGM, both using the above settings. Word accuracy is used as the evaluation metric. Data augmentation like rotation, perspective distortion, motion blur, and gaussian noise, are randomly performed. The maximum text length is set to 25 during training. The size of the character set is set to 94 for English and 6624 [28] for Chinese. In the experiments below, SVTRv2 means SVTRv2-B unless specified. All models are trained on 4 RTX 4090 GPUs.
@@ 4.2 Ablation Study
Effectiveness of MSR. We group Curve and MO text in U14M based on the aspect ratio . As shown in Tab. 1, the majority of irregular texts fall within and , where they are particularly prone to distortion when resized to a fixed size (see Fixed32×128 in Fig. 4). In contrast, MSR demonstrates significant improvements of 15.3% in and 5.2% in compared to Fixed32×128. Meanwhile, a large fixed-size Fixed64×256, although improving the accuracy compared to the baseline, still performs worse than our MSR by clear margins. The results strongly confirm our hypothesis that undesired resizing would hurt the recognition. Our MSR effectively mitigates this issue, providing better visual features thus enhancing the recognition accuracy.
Effectiveness of FRM. We ablate the two rearrangement sub-modules (Horizontal (H) rearranging and Vertical (V) rearranging). As shown in Tab. 1, compared to without FRM (w/o FRM), they individually improve accuracy by 2.03% and 0.71% on MO, and they together result in a 2.46% gain. In addition, we validate the use of a Transformer Block (+ TF1) as an alternative to splitting the process into two steps for learning the matrix holistically. However, its effectiveness is less pronounced, likely because it fails to effectively distinguish between vertical and horizontal orientations. In contrast, FRM performs feature rearrangement in both directions, making it highly sensitive to text irregularity, and thus facilitating accurate CTC alignment. As shown in the left five cases in Fig. 4, FRM successfully recognizes reverse instances, providing strong evidence of FRM’s effectiveness.
@@
| Method | Venue | Encoder | Common Benchmarks (Com) | Avg | Union14M-Benchmark (U14M) | Avg | LTB | OST | Size | FPS | ||||||||||||||||||||||||||
@@
| CPPD [13] | TPAMI25 | SVTR-B | 99.0 | 97.8 | 98.2 | 90.4 | 94.0 | 99.0 | 96.40 | 86.2 | 78.7 | 76.5 | 82.9 | 83.5 | 81.9 | 83.5 | 81.91 | 0.0 | 79.6 | 27.0 | 125 | |||||||||||||||
| IGTR-AR [14] | TPAMI25 | SVTR-B | 98.7 | 98.4 | 98.1 | 90.5 | 94.9 | 98.3 | 96.48 | 90.4 | 91.2 | 77.0 | 82.4 | 84.7 | 84.0 | 84.4 | 84.86 | 0.0 | 76.3 | 24.1 | 58.3 | |||||||||||||||
| SMTR [12] | AAAI25 | FocalSVTR | 99.0 | 97.4 | 98.3 | 90.1 | 92.7 | 97.9 | 95.90 | 89.1 | 87.7 | 76.8 | 83.9 | 84.6 | 89.3 | 83.7 | 85.00 | 55.5 | 73.5 | 15.8 | 66.2 | |||||||||||||||
@@ 4.3 Comparison with State-of-the-arts
We compare SVTRv2 with 24 popular STR methods on Com, U14M, OST, and LTB. The results are presented in Tab. 3. SVTRv2-B achieves top results in 9 out of the 15 evaluated scenarios and outperforms the most of EDTRs, showing a clear accuracy advantage. Meanwhile, it enjoys a small model size and a significant speed advantage. Specifically, compared to MAERec, the best-performed existing model on U14M, SVTRv2-B shows an accuracy improvement of 0.97% and 8 faster inference speed. Compared to CPPD, which is known for its wonderful accuracy-speed tradeoff, SVTRv2-B runs faster than 10%, along with a 4.23% accuracy increase on U14M. Regarding OST, as illustrated in the right part of Fig. 4, SVTRv2-B relies solely on a single visual model but achieves comparable accuracy to PARSeq, which employed the advanced permuted language model and is the best-performed existing model on OST. In addition, SVTRv2-T and SVTRv2-S, the two smaller models also show leading accuracy compared with models of similar sizes, offering flexible solutions with different accuracy-speed tradeoff.
Two observations are derived by looking at the results on Curve and MO. First, SVTRv2 models significantly surpass existing CTC models. For example, compared to SVTR-B, SVTRv2-B gains prominent accuracy improvements of 14.4% and 44.5% on the two subsets, respectively. Second, as shown in Tab. 4, comparing to previous methods employing rectification modules [40, 36, 59, 65, 54, 11, 64] or attention-based decoders [38, 5, 51, 56, 25, 54, 52, 46, 29] to recognize irregular text, SVTRv2 also performs better than these methods on Curve. In Fig. 4, TPS (a rectification module) and MAERec* (an attention-based decoder) do not recognize the extremely curved and rotated text correctly. In contrast, SVTRv2 successes. Moreover, as demonstrated by the results on LTB in Tab. 4 and Fig. 5, TPS and MAERec* both do not effectively recognize long text, while SVTRv2 circumvents this limitation. These results indicate that our proposed modules successfully address the challenge of handling irregular text that existing CTC models encountered, while still preserving CTC’s proficiency in recognizing long text.
SVTRv2 also exhibit strong performance in Chinese text recognition (see Tab. 5), where SVTRv2-B achieve state of the art. The result implies its great adaptability to different languages. Moreover, it also shows superior performance on Chinese long text (SceneL>25). To sum, we evaluate SVTRv2 across a wide range of scenarios. The results consistently confirm that this CTC model beats leading EDTRs.
| Curve | MO | Com | U14M | LTB | ||||||
@@
| Method | Common Benchmarks (Com) | Avg | OST | Size | FPS | |||||
| E2STR [63] | 99.2 | 98.6 | 98.7 | 93.8 | 96.7 | 99.3 | 97.71 | 80.7 | 211 | 7.86 |
| VL-Reader [66] | 99.6 | 99.1 | 98.7 | 92.6 | 97.5 | 99.3 | 97.80 | 86.2 | 142 | - |
| CLIP4STR [62] | 99.4 | 98.6 | 98.3 | 90.8 | 97.8 | 99.0 | 97.32 | 82.8 | 158 | 14.1 |
| DPTR [61] | 99.5 | 99.2 | 98.5 | 91.8 | 97.1 | 98.6 | 97.45 | - | 66.5 | 49.3 |
| IGTR [14] | 99.2 | 98.3 | 98.8 | 92.0 | 96.8 | 99.0 | 97.34 | 86.5 | 24.1 | 58.3 |
| SVTRv2-B | 99.2 | 98.6 | 98.8 | 93.8 | 97.2 | 99.4 | 97.83 | 86.9 | 19.8 | 143 |
In addition, recent EDTRs advances, e.g., E2STR [63], VL-Reader [66], CLIP4STR [62], and DPTR [61], achieve impressive accuracy through large-scale vision-language pretraining. To align with these methods, we conduct an experiment by adding pretraining to SVTRv2 on synthetic datasets [20, 24] and fine-tuning on Real dataset [4]. The results in Tab. 6 show that this pretraining significantly enhances SVTRv2’s performance, allowing it to surpass the aforementioned models. Notably, SVTRv2 achieves the highest average accuracy in Com (97.8%) while also demonstrating superior generalization to OST (86.9%). Compared to CLIP4STR, SVTRv2 achieves these results with only 14% of the parameters and runs 10 faster, highlighting its efficiency. These findings again validate the effectiveness of our SVTRv2, as well as the proposed strategies or modules, i.e., MSR, FRM, and SGM.
@@ 5 Conclusion
In this paper, we have presented SVTRv2, an accurate and efficient CTC-based STR method. SVTRv2 is featured by developing the MSR and FRM modules to tackle the text irregular challenge, and devising the SGM module to endow linguistic context to the visual model. These upgrades maintain the simple inference architecture of CTC models, thus they remain quite efficient. More importantly, our thorough validation on multiple benchmarks demonstrates the effectiveness of SVTRv2. It achieves leading accuracy in various challenging scenarios covering regular, irregular, occluded, Chinese and long text, as well as whether employing pretraining. In addition, we retrain 24 methods from scratch on U14M-Filter without data leakage. Their results on U14M constitutes a comprehensive and reliable benchmark. We hope that SVTRv2 and this benchmark will further advance the development of the OCR community.
Acknowledgement This work was supported by the National Natural Science Foundation of China (Nos. 62427819, 32341012, 62172103).
@@
| ID | Method | Common Benchmarks (Com) | Avg | Union14M-Benchmark (U14M) | Avg | LTB | OST | Size | FPS | |||||||||||||||||||||||||
@@
6 More Details of Ablation Study
SVTRv2 builds upon the foundation of SVTR by introducing several innovative strategies aimed at addressing challenges in recognizing irregular text and modeling linguistic context. The key advancements and their impact are detailed as follows:
Removal of the rectification Module and introduction of MSR and FRM. In the original SVTR, a rectification module is employed to recognize irregular text. However, this approach negatively impacts the recognition of long text. To overcome this limitation, SVTRv2 removes the rectification module entirely. To effectively handle irregular text without compromising the CTC model’s ability to generalize to long text, MSR and FRM are introduced.
@@ 6 More Details of Ablation Study
Refinement of local mixing mechanisms. SVTR employs a hierarchical vision transformer structure, leveraging two mixing strategies: Local Mixing is implemented through a sliding window-based local attention mechanism, and Global Mixing employs the standard global multi-head self-attention mechanism. SVTRv2 retains the hierarchical vision transformer structure and the global multi-head self-attention mechanism for Global Mixing. For Local Mixing, SVTRv2 introduces a pivotal change. Specifically, the sliding window-based local attention is replaced with two consecutive group convolutions (Conv2) [21]. It is important to highlight that unlike previous CNNs, there is no normalization or activation layer between the two convolutions.
Semantic guidance module. The original SVTR model relies solely on the CTC framework for both training and inference. However, CTC is inherently limited in its ability to model linguistic context. SVTRv2 addresses this by introducing a Semantic Guidance Module (SGM) during training. SGM facilitates the visual encoder in capturing linguistic information, enriching the feature representation. Importantly, SGM is discarded during inference, ensuring that the efficiency of CTC-based decoding remains unaffected while still benefiting from its contributions during the training phase.
@@ 6.1 Progressive Ablation Experiments
To comprehensively evaluate the contributions of every SVTRv2 upgrade, a series of progressive ablation experiments are conducted. Tab. 7 outlines the results, along with the following observations:
@@
2. Rectification Module Removal (ID 1) reveals that while the rectification module (e.g., TPS) improves irregular text recognition accuracy, it hinders the model’s ability to recognize long text. This confirms its limitations in balancing different recognition tasks.
@@
4. Replacement of Local Attention with Conv2 (ID 3): Replacing the sliding window-based local attention with two consecutive group convolutions (Conv2) yields improvements in artistic text, with a 3.0% increase in accuracy. This result highlights the efficacy of convolution-based approaches in capturing character-level nuances, such as strokes and textures, thereby improving its ability to recognize artistic and irregular text.
@@
6. Integration of SGM (ID 6): Adding SGM yields significant gains on multiple datasets, improving accuracy on OST by 5.11% and U14M by 2.28%.
It can be summarized as that, by integrating Conv2, MSR, FRM, and SGM, SVTRv2 significantly improves performance in recognizing irregular text and modeling linguistic context over SVTR, while still maintaining robust long-text recognition capabilities and preserving the efficiency of CTC-based inference.
@@ 7 SVTRv2 Variants
In Tab. 8, denotes that the first m mixing blocks in SVTRv2 utilize local mixing, while the last n mixing blocks employ global mixing. Specifically, in SVTRv2-T and SVTRv2-S, all blocks in the first stage and the first three blocks in the second stage use local mixing. The last three blocks in the second stage, as well as all blocks in the third stage, are global mixing. In the case of SVTRv2-B, all blocks in the first stage and the first two blocks in the second stage use local mixing, whereas the last four blocks in the second stage and all blocks in the third stage adopt global mixing.
@@
| Curve | Multi-Oriented | Artistic | Contextless | Salient | Multi-Words | General | |
@@
8 More Details of Real-World Datasets
For English recognition, we train models on real-world datasets, from which the models exhibit stronger recognition capability [4, 25, 37]. There are three large-scale real-world training sets, i.e., the Real dataset [4], REBU-Syn [37], and Union14M-L (U14M-Train) [25]. However, as shown in Fig. 6 and Tab. 9, the former two significantly overlap with U14M, thus not suitable for model training when using U14M at the evaluation dataset. Surprisingly, U14M-Train is also overlapped with U14M in nearly 6.5k text instances across the seven subsets. It means the models trained based on U14M-Train suffer from data leakage when tested on U14M, thus the results reported by [25] should be updated. To this end, we create a filtered version of Union14M-L, termed as U14M-Filter, by filtering these overlapping instances from the training set. This new dataset is used to train SVTRv2 and other 24 methods we reproduced.
9 More Details of Inference Time
In terms of the inference time, we do not utilize any acceleration framework and instead employ PyTorch’s dynamic graph mode on one NVIDIA 1080Ti GPU. We first measure the inference time for 3,000 images with a batch size of 1, calculating the average inference time for each text length. We then compute the arithmetic mean of the average time across all text lengths to determine the overall inference time of the model. Algorithm 1 details the process of measuring inference time.
10 Results when Trained on Synthetic Datasets
Previous research typically follows a typical evaluation protocol, where models are trained on synthetic datasets and validated using Com, the six widely recognized real-world benchmarks. Following this protocol, we also train SVTRv2 and other models on synthetic datasets. In addition to evaluating SVTRv2 on Com, we assess its performance on U14M. The results offer a comprehensive evaluation of the model’s generalization capabilities. For methods that have not reported performance on challenging benchmarks, we conduct additional evaluations using their publicly available models and present these results for comparative analysis. As illustrated in Tab. 10, models trained on synthetic datasets exhibit notably lower performance compared to those trained on large-scale real-world datasets (see Tab. 3). This performance drop is particularly pronounced on challenging benchmarks. These findings highlight the critical importance of real-world datasets in improving recognition accuracy.
Despite trained on less diverse synthetic datasets, SVTRv2 also exhibits competitive performance. On irregular text benchmarks, such as Curve and Multi-Oriented, SVTR achieves strong results, largely due to its integrated rectification module [40], which is particularly adept at handling irregular text patterns, even when trained on synthetic datasets. Notably, SVTRv2 achieves a substantial 4.8% improvement over SVTR on Curve, further demonstrating its enhanced capacity to address irregular text. Overall, these results demonstrate that, even when trained solely on synthetic datasets, SVTRv2 exhibits strong generalization capabilities, effectively handling complex and challenging text recognition scenarios.
| Method | Venue | Encoder | Common Benchmarks (Com) | Avg | Union14M-Benchmark (U14M) | Avg | Size | |||||||||||||||||||||||||